OpenStudy (anonymous):

I award medals to whoever helps me!!! given f(x)=3x-1 over 2. Solve for f^-1 (4)

OpenStudy (anonymous):

1 3 5 7

OpenStudy (anonymous):

wow this is tough

OpenStudy (anonymous):

lol tell me about it! I have been working on this single problem for 30 minutes trying to figure out how to solve it

OpenStudy (anonymous):

and I have had 0 progress with it :p

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i've got nothing sry man

OpenStudy (anonymous):

it's okay :-) Thanks for trying!

OpenStudy (anonymous):

I gave you a medal because you tried to help me :-) Nobody else has and I posted this on another thing too

OpenStudy (anonymous):

nah dude keep the medal i don't deserve it

OpenStudy (phi):

f−1(x) means the inverse of f(x) it is asking for the x that makes f(x)= 4 we can find f−1(x) (though this is not necessary) start with y = (3x-1)/2 swap x and y x = (3y -1)/2 solve for y 2x = 3y -1 2x + 1 = 3y y = (2x+1)/3 now plug in 4 for x to get y = (2*4+1)/3 = 9/3 = 3

OpenStudy (phi):

another way is to solve for x in f(x)=4 4 = (3x-1)/2 multiply both sides by 2 8 = 3x -1 add 1 to both sides 9 = 3x divide both sides by 3 3 = x

OpenStudy (anonymous):

Oh wow!! That clears everything up so much! Thank you SO much!!

OpenStudy (anonymous):

f(x)=4 4 = (3x-1)/2 X both sides 2 8 = 3x -1 +1 to both sides 9 = 3x / 3 x=3 , Hope it helped! Could use a medal if it did!

OpenStudy (anonymous):

Thanks Hannah! It did help but I already gave the medal to phi. You both helped me a lot! If it helps, I became your fan and your smart score was boosted from 19 to 24.